OpenStudy (anonymous):

please help!!! Students at Smith Middle School submitted their votes to name the school's snake mascot. The results from one of the ballot boxes are shown in the table. Name the Mascot Name Votes SSSneaky 8 SSSteve 7 SSSammy 4 SSStanley 6 SSStriker 15 If two ballots are chosen randomly one at a time from the ballot box, without replacing the first one, what is the probability that both ballots are votes for SSSammy? i think it might be B or D A. 1/130 B. 2/195 C. 1/100 D. 3/400

OpenStudy (anonymous):

how many votes are all together?

OpenStudy (anonymous):

41

OpenStudy (anonymous):

So why are the answers : 1/ (100) and so on?

OpenStudy (anonymous):

There is a vote missing then? 8+7+4+6+15 = 40

OpenStudy (anonymous):

something is definitely missing here

OpenStudy (anonymous):

Anyway, assuming that there are 40 votes, we have 4/40 chances to get the first ballot as SSSamy AND 3/39 to get the second ballot as SSSamy. (1/10)(3/39) = 3/390 = 1/130.

OpenStudy (anonymous):

If you have a 0.01 percentage, than the closest to that it 1/100

OpenStudy (anonymous):

But it's not a 1% chance, it's less than that, it's 3/390 = 0.77%, approx.

OpenStudy (anonymous):

But, if you were to relate the probability respectively from 4/40 and 1/100, would that not be equivalent?

OpenStudy (anonymous):

Yeah, that's the chance for one of the ballots to be SSSamy. Then, after we take one SSSamy, there are 3 ballots left and 39 total. "without replacing the first one" is the key here.

OpenStudy (anonymous):

Which is equivalent to 1/13.

OpenStudy (anonymous):

Okay, but is that neccessary? I thought we only needed Ssamy's probability? I'm probably wrong here :o but thanks for putting up with it

OpenStudy (anonymous):

"both ballots are votes for SSSammy?" We want two ballots in a row to be votes for SSSamy.

OpenStudy (anonymous):

Ah. I see now, you're correct.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

No problem :-)